y = 3e 2x + 2e 3x ..prove that d 2 y/dx 2 – 5dy/dx + 6y =0

y = 3e2x + 2e3x..prove that d2y/dx2 –  5dy/dx + 6y =0

Grade:12

1 Answers

Adarsh
6 years ago
Dy/DX=6e^2x+6e^3xD2y/dx2=12 e^2x+18e^3xSubstituting in the equation we get12e^2x+18e^2x-30e^2x+18e^3x+12e^3x-30e^3x=30e^2x-30e^2x+30e^3x-30e^3x=0.Hence proved